OpenStudy (anonymous):

If you save three pennies on January 1, six pennies on January 2, nine pennies on January 3, and continue this pattern for one year (not a leap year), what will be the value of your entire savings, in dollars, at the end of that one year? Express your answer as a decimal. ***Please show steps and explain!! :)

OpenStudy (amistre64):

it sounds like the pattern is the 3s times table: 3,6,9,12,15,...

OpenStudy (amistre64):

but we have to add that amount to the previous balance

OpenStudy (amistre64):

it develops into a recursive equation b1 = 3 b2 = 3 + 6 = 9 b3 = 3 + 6 + 9 = 18 b4 = 3 + 6 + 9 + 12 = 40 bn = b{n-1} + 3(n-1) maybe

OpenStudy (amistre64):

from my memory banks; i think this might end up with some exponential closed form

OpenStudy (amistre64):

\[\sum_{n=1}^{365}3n\] perhaps is another way to notate it

OpenStudy (anonymous):

i have no idea how to solve this :( so the equation we would use would be An=3+3(n-1) ??

OpenStudy (amistre64):

we can test out the 3+3(n-1) and see how it might play out when n=1; a=3 when n=2; a=6 when n=3; a=9 that works for the amount of money you put in each time, but it doesnt keep track of the balance that accumulates

OpenStudy (anonymous):

Oooops! I meant this: \[\large \sum_{k=1}^nk = \frac{n(n+1)}{2}\]

OpenStudy (anonymous):

wait so do i use wio's equation to solve?? i'm confused... :O cuz i need to use an equation right?? otherwise that would take forever to plug in numbers and stuff??

hartnn (hartnn):

if u find sum of the series 3+3(n-1) till 365 terms.....u can get it.....maybe..

OpenStudy (amistre64):

wios formula is for the sum of an arithmetic series with "k" terms im not sure if its the proper setup for you problem tho; but it does play a part

OpenStudy (anonymous):

ohhh okat so what equation do i use then?

OpenStudy (amistre64):

3(1) + 3(2) + 3(3) + ... + 3(365) 3(1+2+3+ ... +365)

OpenStudy (anonymous):

okay, following :)

OpenStudy (amistre64):

\[3\frac{1+365}{2}\]should do it

hartnn (hartnn):

3*n(n+1)/2 n=365

OpenStudy (anonymous):

could i use this equation?? Sn=(n/2)[2a+(n-1)d]

hartnn (hartnn):

yes, same answer u will get

OpenStudy (anonymous):

oh so n=365 represents number of terms??

OpenStudy (amistre64):

... i forgot the extra "n" in the latex :/

OpenStudy (anonymous):

You mean \[\Large 3 \cdot \frac{365(365+1)}{2}\]

OpenStudy (amistre64):

correct wio ;)

hartnn (hartnn):

yup, n= no. of terms = 365 here and if this question belongs to topic series, the u use that sn formula only....

OpenStudy (anonymous):

So if i plug into your equation who, then i get 200385 and if i plug into my equation, i get this: n= 365 terms a=1st term which = 3 d= common difference b/w each terms which is 3 Sn=(365/2)[2*3+(365-1)3] =(365/2)[6+(364)3] =(365/2)[6+1092] =(365/2)(1098) =365*549 =200385 ????

OpenStudy (anonymous):

is that right??? cuz i got the same thing twice???

hartnn (hartnn):

yup, even i got that

OpenStudy (amistre64):

lol, with any luck when you get the same thing twice its correct; but theres always the possibility that you might of just done 2 things incorrectly

hartnn (hartnn):

these are pennies, u need to convert them in dollars

OpenStudy (anonymous):

and so the simplest way to show the work would probably be